solve the system of first-order linear differential equations

y1'=y1-y2

y'2=2y1+4y2


1
Expert's answer
2021-10-27T13:57:08-0400

Let us solve the system of first-order linear differential equations

"\begin{cases}
y_1'=y_1-y_2\\
y_2'=2y_1+4y_2
\end{cases}"

It follows that "y_2=y_1-y_1'," and hence "y_2'=y_1'-y_1''." Consequently, we get the equation

"y_1'-y_1''=2y_1+4(y_1-y_1')," which is equivalent to "y_1''-5y_1'+6y_1=0."


The characteristic equation "k^2-5k+6=0" of "y_1''-5y_1'+6y_1=0" is equivalent to "(k-2)(k-3)=0," and thus it has the roots "k_1=2,\ k_2=3." Therefore the solution of this differential equation is "y_1=C_1e^{2x}+C_2e^{3x}."

Then "y_1'=2C_1e^{2x}+3C_2e^{3x}," and therefore

"y_2=y_1-y_1'=C_1e^{2x}+C_2e^{3x}-(2C_1e^{2x}+3C_2e^{3x})=-C_1e^{2x}-2C_2e^{3x}."


We conclude that the general solution of the system is of the form:

"\begin{cases}
y_1=C_1e^{2x}+C_2e^{3x}\\
y_2=-C_1e^{2x}-2C_2e^{3x}
\end{cases}."